AnimeGhoul8863:

Which graph best represents the function f(x) = 1(1.5)x? graph of increasing exponential function going through point 0, 2 graph of increasing exponential function going through point 0, 3 graph of increasing exponential function going through point 0, 1 graph of increasing exponential function going through point 0, 4

Ultrilliam:

@563blackghost

Ultrilliam:

Welcome to QC, I tagged someone who should be able to help :)

AnimeGhoul8863:

ok thank you

Nnesha:

you can replace x and y with given points if both sides of the equation equal then that point will be the answer

AnimeGhoul8863:

so its B?

Nnesha:

\[\large\rm (\color{Red}{x} ,\color{blue}{y} ) \rightarrow (\color{red}{0},\color{blue}{2})\] \[\large\rm \color{blue}{f(x)}=1(1.5)^\color{red}{x}\]\[\large\rm \color{blue}{2}=1(1.5)^\color{red}{0}\] both sides are equal ??

Nnesha:

how did u get B ?

AnimeGhoul8863:

Im sooooo sorry i miss read it so its A

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha \[\large\rm (\color{Red}{x} ,\color{blue}{y} ) \rightarrow (\color{red}{0},\color{blue}{2})\] \[\large\rm \color{blue}{f(x)}=1(1.5)^\color{red}{x}\]\[\large\rm \color{blue}{2}=1(1.5)^\color{red}{0}\] both sides are equal ?? \(\color{#0cbb34}{\text{End of Quote}}\) sole the right side .. is it equal to the left side ???

Nnesha:

solve*

AnimeGhoul8863:

Umm.... if im correct the answer is D

AnimeGhoul8863:

and sorry for causing trouble im just really not good at math

Nnesha:

the problem is not that u r good or bad when it comes to math the problem is you're not trying just guessing the answer

AnimeGhoul8863:

so its not D how i dont understand

Nnesha:

how it can be D that's what i don't get it. Btw you haven't answer my question

Ultrilliam:

Lemme see if I can simplify what she is saying.

AnimeGhoul8863:

what was ur question

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha \[\large\rm (\color{Red}{x} ,\color{blue}{y} ) \rightarrow (\color{red}{0},\color{blue}{2})\] \[\large\rm \color{blue}{f(x)}=1(1.5)^\color{red}{x}\]\[\large\rm \color{blue}{2}=1(1.5)^\color{red}{0}\] both sides are equal ?? \(\color{#0cbb34}{\text{End of Quote}}\) 1(1.5)^0 is equal to what ?? is it equal to 2 ?

Ultrilliam:

Essentially you have the left and right part of the problem for each answer choice. 1*1.5^the left side of the problem 1*1.5^0=x After you find that what your going to want to do is see if the value of that is equal to any of the other right side of the problems. That said, what does 1*1.5^0 equal?

AnimeGhoul8863:

1*1.5=1.5

Ultrilliam:

and what is 1.5^0?

Ultrilliam:

(HINT: Any number to the power of 0 is 1!)

AnimeGhoul8863:

1.5^0=1 so the answer would be 0,1 wow that was confusing but i get it

AnimeGhoul8863:

no ones been able to get me to understand math before but you just did so thanks

Ultrilliam:

Your very welcome :)

AnimeGhoul8863:

so who do i give best response to

Ultrilliam:

Anyone you want, I wouldn't have been able to help you without @Nnesha though =P

AnimeGhoul8863:

is it ok to post another question

Ultrilliam:

Yep, you'll have to close your current question first though

AnimeGhoul8863:

ok thx

Eiwoh2:

This was apparently your first question? XD